Before working this problem, review Conceptual Example 14. A pellet gun is fired straight downward from the edge of a cliff that is 15.5 m above the ground. The pellet strikes the ground with a speed of 34.9 m/s. How far above the cliff edge would the pellet have gone had the gun been fired straight upward? (Please write out on paper)
Answer #1

first we need to find velocity with which bullet projected

using 3rd equation of motion

v^2 = u^2 + 2 g h

34.9^ = u^2 + 2*9.8* 15.5

u = 30.236 m/s

now, when bullet projected vertically upward

0 = u^2 - 2 g H

30.236^2 = 2*9.8 * H

H = 46.643 m

===========

do comment in case any doubt, will reply for sure.. Goodluck